OpenStudy (ar43r):

find x: 1/2*ln(15/20)=ln(1-x)

OpenStudy (anonymous):

1/2 ln(15/20) =(1/2) ln(3/4) = ln( (3/4)^(1/2)) so ln( (3/4)^(1/2)) = ln(1 - x) (3/4)^(1/2) = 1 - x x = 1 - (3/4)^(1/2)